Solve using the principle of zero products. thank you
(21 - 0.8x)(5 - 0.4x) = 0

Answer by stanbon(75887)

Solve using the principle of zero products. thank you
(21 - 0.8x)(5 - 0.4x) = 0
---
Since the product is zero, at least one of those factors
must be zero.
----------------
If 21-0.8x = 0
then 0.8x = 21
and x = 26.25
-----------------------
OR
---
If 5-0.4x = 0
then 0.4x = 5
and x = 12.5
==========================
Cheers,
Stan H.